OpenStudy (anonymous):

Medal for whoever helps with both questions. 1.Match each star with the letter it goes with. A.Transitive Property of Equality B.Closure of Real Numbers C.Reflexive Property of Equality D.Symmetric Property of Equality *3/4 = 3/4 *If a + 9 = 20, then 20 = a + 9. *If 10 = 4x and 4x = c, then 10 = c. *-18 + 4 is a unique real number -14. 4.Match each star with the letter it goes with. A.Additive Inverse Property B.Multiplicative Identity Property C.Multiplicative Inverse Property D.Additive Identity Property E.Multiplicative Property of 0 *2/3 • 0 = 0 *-3/4 • -4/3 = 1 *1 • 4.5 = 4.5 *-8.9

OpenStudy (anonymous):

3/4=3/4: (reflexive) a+9=20 implies 20=a+9 (symmetric) 10=4x and 4x=c implies 10=c ( transitive) -18+4 is the uniquer real number -14 (closure)

OpenStudy (anonymous):

2/3*0=0 (E) -3/4*-4/3=1 (C) 1*4.5=4.5 (B) *-8.9 ? is there supposed to be more to this?

OpenStudy (anonymous):

*-8.9 + 8.9 = 0 *-8.9 + 0 = -8.9

OpenStudy (anonymous):

-8.9+8.9=0 is additive inverse -8.9+0=-8.9 is additive identity

OpenStudy (anonymous):

do you have questions on these?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

#1 D.Symmetric Property of Equality C.Reflexive Property of Equality A.Transitive Property of Equality B.Closure of Real Numbers